OpenStudy (anonymous):

Find the domain of 4x+3?

OpenStudy (aroub):

4x+3=0 4x=-3 x=-3/4

OpenStudy (anonymous):

x=-3/4

OpenStudy (anonymous):

The domain of 4x + 3 is ALL real numbers. You can plug in any number safely without getting any problems.

OpenStudy (anonymous):

And aroub/nancy: even if you were solving for x, we don't have any assurance that it's 4x+3 = 0, so you can't just assume.

OpenStudy (anonymous):

Nancy Lam and aroub ignoring the prompt lol

OpenStudy (anonymous):

its all real like uber said btw

OpenStudy (anonymous):

Domain, by the way, means "what x values can I plug in without causing any problems?" For example: \[\sqrt{?}\] 's domain would be everything from 0 upwards to infinity. Anything below 0 would have a fake answer.

OpenStudy (saifoo.khan):

-3/4

OpenStudy (anonymous):

lol. Another one.

OpenStudy (aroub):

oh then u can put any number for x ??

OpenStudy (anonymous):

oh okay. Thanks

OpenStudy (anonymous):

Yup. Since it doesn't say 3x+4 = 0, we have to assume it's 3x+4 = y Which means we can put anything in there to get any real output.

OpenStudy (aroub):

oh yea right! thank you! :D

OpenStudy (anonymous):

Not a problem.